KyledaGreat:

The value of a machine, V, at the end of t years is given by V=C(1−r)t, where C is the original cost and r is the rate of depreciation. Find the value of a machine at the end of 2 years if the original cost was $751 and r=0.3. Round to the nearest cent.

Vocaloid:

similar logic as before, just with subtraction since we're depreciating not gaining value V=C(1−r)t = 751(1 - 0.3)^2

KyledaGreat:

the answer is 367.99

Vocaloid:

good

KyledaGreat:

Thank you